# Topology and Algebra of Bonded Knots and Braids

### v1

@@ 9.2. The bonded semi-closures of bonded knotoids
Obviously, different closures of a bonded knotoid may result in non isotopic bonded knots. Thus, in order to represent bonded knots via bonded knotoid diagrams, we fix the closure type and we have the following result (see also [31] for the case of knotoids):
@@
Assuming a specific closure type, there is a well-defined surjective map from bonded knotoids to bonded knots.
@@
In [27] the authors introduce the notion of singular knotoids, that is, knotoid diagrams that contain a finite number of self-intersections, and they present a different type of closure operation for knotoids, the singular closure. In analogy to the definition of singular closure (under certain conditions) in [6] the pseudo closure of knotoids is defined, where a pseudo knotoid diagram is a knotoid with some missing crossing information.
In the theory of bonded knotoids, a different “closure” operation may be considered. In particular, if we allow the nodes of a bond to coincide with the endpoints of a knotoid, we obtain the bonded semi-closure of a bonded knotoid. The term semi here denotes the fact that this operation does not yield a bonded knot. It can be interpreted as an indication that the two endpoints of the knotoid are attracted by a physical force. More precisely, we have the following definition:
@@
We call the bonded knotoid obtained by connecting the endpoints of a bonded knotoid diagram with a bond that goes under or over each arc and each bond it meets, the underpass semi-closure and the overpass semi-closure of the bonded knotoid respectively (see Figure 78).
Figure 78. The bonded semi-closures of a bonded knotoid.
We believe that this closure operation, interpreted as a physical force between the endpoints of a knotoid diagram, could play an important role in the study of open proteins.
@@
A braidoid diagram $B$ is a system of a finite number of arcs immersed in $[0,1]\times[0,1]\subset\mathbb{R}^{2}$, where $\mathbb{R}^{2}$ is identified with the xt-plane, such that the t-axis is directed downward. The arcs of $B$ are the strands of $B$. Each strand is naturally oriented downward, with no local maxima or minima, following the natural orientation of $[0,1]$. Moreover, there are only finitely many intersection points among the strands, which are transversal double points endowed with over/under data, and are called crossings of $B$. A braidoid diagram has two types of strands, the classical strands, i.e. braid strands connecting points on $[0,1]\times\{0\}$ to points on $[0,1]\times\{1\}$, and the free strands that either connect a point in $[0,1]\times\{0\}$ or in $[0,1]\times\{1\}$ to an endpoint located anywhere in $[0,1]\times[0,1]$, or they connect two endpoints that are located anywhere in $[0,1]\times[0,1]$. These points that lie on $[0,1]\times\{0\}$ or $[0,1]\times\{1\}$, are called braidoid ends.
@@
    Bonded swing moves as illustrated in Figure 80: the endpoints are allowed to swing to the right or the left like a pendulum as long as the downward orientation on the moving arc is preserved, and the forbidden moves are not violated.
@@
We first treat strands of the braids that contain the nodes of bonds as shown in Figures 57 and 58. In that way we obtain up-arcs that do not contain the nodes of bonds. We apply then the braidoiding algorithm of [12] (which is similar to the braidoiding algorithm we presented in § 6 and we obtain a bonded braidoid, whose closure is isotopic to the bonded multi-knotoid. ∎
@@ 10. Further Work
Along similar lines, we will formulate $L$-equivalence in the setting of bonded braidoids, extending the braidoid – knotoid and the bonded link – bonded braid correspondences to the bonded and enhanced bonded braidoid categories.
Along with the braid category, one can also formulate a Bonded Morse Category for multi-knotoids and linkoids. Bonds are included just as with braids, but in the Morse category maxima (caps) and minima (cups) are allowed in the diagrams. Thus there are generators for a braided tensor category consisting in local bonds, crossings of two types, cups and caps and identity lines. There is an extension of Reidemeister moves for this Morse category. Then one has representation theorems that assure that entities can be written in Morse form, and generalized Reidemeister moves imply that topologically equivalent entities can be connected by sequences of moves. This categorical approach allows the construction of invariants of bonded Morse diagrams in terms of functors from the Morse category to a category of modules and traces. For a formulation of a Morse category for knotoids see [14]. We will detail this approach in a subsequent paper.

### v2

@@ 9.2. Types of closures for bonded multi-knotoids
Obviously, different closures of a bonded multi-knotoid may result in non isotopic bonded links. On the other hand, we can represent bonded links via bonded multi-knotoid diagrams, by cutting any arc and declaring the two ends as endpoints. If we fix the closure type we have the following result (proved as in [41] for the case of knotoids):
@@
Assuming a specific closure type, there is a well-defined surjective map from planar (or spherical) bonded multi-knotoids to bonded links.
@@
In [34] the authors introduce the notion of singular knotoids, that is, knotoid diagrams that contain also singular crossings, and they present a different type of closure operation for singular knotoids, the singular closure, whereby the closing arcs crosses singularly every arc it intersects. In analogy to the definition of singular closure (under certain conditions) in [9] the pseudo closure of knotoids is defined, where a pseudo knotoid diagram is a knotoid with some missing crossing information.
Likewise, in the theory of bonded multi-knotoids, a different “closure” operation may be considered, whereby we introduce a bond whose nodes coincide with the endpoints of a bonded multi-knotoid. More precisely, we have the following definition:
@@
The operation whereby the endpoints of a bonded multi-knotoid diagram are connected with a standard bond shall be called bonded closure. In particular, if the connecting bond passes under resp. over each arc and each bond it meets, then we have the underpass bonded closure, resp. the overpass bonded closure of the bonded multi-knotoid (see Figure 84).
Figure 84. Bonded closures of a bonded knotoid.
Note that the bonded closure does not yield a bonded link. The bonded closure can be interpreted as an indication that the two endpoints of the multi-knotoid are attracted by a physical force. We believe that this closure operation with the above interpretation could play an important role in the study of open proteins.
The theory of enhanced bonded links carries through in complete analogy to the bonded multi-knotoids, to obtain the theory of enhanced bonded multi-knotoids. Only, for the bonded closure it rather makes sense to consider only the bonded closure with an attracting bond.
@@
[…]
@@
    Bonded swing moves as illustrated in Figure 86: the endpoints are allowed to swing to the right or the left like a pendulum as long as the downward orientation on the moving arc is preserved, and the forbidden moves are not violated. A swing move could include a bonded arc or not, see Figure 86.
@@
We first treat strands of the braids that contain the nodes of bonds as shown in Figures 62 and 63. In that way we obtain up-arcs that do not contain the nodes of bonds. We apply then the braidoiding algorithm of [18] (which is similar to the braidoiding algorithm we presented in § 6 and we obtain a bonded braidoid, whose closure is isotopic to the bonded multi-knotoid. ∎
@@ 10. Further Work
We will formulate $L$-equivalence in the setting of bonded braidoids, extending the correspondences to the bonded and enhanced bonded braidoid categories. We will define the plat closure for bonded braids and braidoids and will formulate Hilden and Birman type theorems [21, 6].
We will explore a bonded Morse category for multi-knotoids and linkoids. Bonds are included just as with braids, but in the Morse category maxima (caps) and minima (cups) are allowed in the diagrams. Thus there are generators for a braided tensor category consisting in local bonds, crossings of two types, cups and caps and identity lines. There is an extension of Reidemeister moves for this Morse category. Then one has representation theorems that assure that entities can be written in Morse form, and generalized Reidemeister moves imply that topologically equivalent entities can be connected by sequences of moves. This categorical approach allows the construction of invariants of bonded Morse diagrams in terms of functors from the Morse category to a category of modules and traces. For a formulation of a Morse category for knotoids see [17]. We will detail this approach in a subsequent paper.